# Grothendieck’s proof of Hirzebruch-Riemann-Roch theorem

### v1

@@ 2.1 The Grothendieck group
Example 2.6.
Let $X$ be a projective variety over a field $k$, then we have a well defined homomorphism
$\chi\left(X,\bullet\right):G_{0}\left(X\right)$ $\to$ $\mathbb{Z}$
$\left[\mathcal{F}\right]$ $\mapsto$ $\sum_{i\geq 0}\left(-1\right)^{i}\dim_{k}H^{i}\left(X,\mathcal{F}\right)$
We see that the sum is actually finite in view, for example, of Grothendieck’s Vanishing [Har, Theorem III.2.7, page 208] and each term is finite-dimensional in view of [Har, Theorem III.5.2, page 228]. In order to see that this defines a group homomorphism, consider a short exact sequence in $\mathrm{Coh}_{X}$
$0\to\mathcal{F}\to\mathcal{G}\to\mathcal{H}\to 0$
and the associated long exact sequence of finite-dimensional $k$-vector spaces
$0\to H^{0}\left(X,\mathcal{F}\right)\to H^{0}\left(X,\mathcal{G}\right)\to H^{0}\left(X,\mathcal{H}\right)\to\dots\to H^{n}\left(X,\mathcal{H}\right)\to 0$
with $n=\dim X$. By additivity of the dimension on exact sequences we conclude that
$\chi\left(X,\mathcal{G}\right)=\chi\left(X,\mathcal{F}\right)+\chi\left(X,\mathcal{H}\right).$
Of course, exactly the same line shows that $\chi\left(X,\bullet\right)$ is defined as $K_{0}\left(X\right)\to\mathbb{Z}$.
@@
Let $\mathcal{P}$ be a category with exact sequences seen as a full subcategory of an abelian category $\mathcal{C}$ and suppose that, for every epimorphism $\pi:X\to Y$ in $\mathcal{P}$, $\ker\left(\pi\right)$ is still an object of $\mathcal{P}$. Let
@@
be an eact sequence in $\mathcal{P}$. Then the equality
@@
Let us proceed by induction on $n$: the cases when $n\leq 2$ are trivial. Suppose the statement holds forn1n-1, then we can consider the two exact sequences
@@
Proposition 2.9.
Let $A$ be a regular ring [^6] We say that a local ring $\left(A,\mathfrak{m},k\right)$ is a regular local ring if it is noetherian and $\dim A=\dim_{k}\frac{\mathfrak{m}}{\mathfrak{m}^{2}}$ We say that a (non necessarily local) ring $A$ is regular if it is noetherian and for every prime ideal $\mathfrak{p}\in\mathrm{Spec}\left(A\right)$, the local ring $A_{\mathfrak{p}}$ is regular. (resp. let $X$ be a smooth projective variety). Then the natural inclusions $K_{0}\left(A\right)\to G_{0}\left(A\right)$ and $K_{0}\left(X\right)\to G_{0}\left(X\right)$ are isomorphisms. [^7] Look at the Resolution Theorem [Ros, Theorem 3.1.13, pag. 121] (or the subsequent page for the $K_{1}$-analogue) for a generalization of this result. In general one has a pletora of dévissage theorems for reducing the study of a category to that of a well-behaved subcategory by means of suitable exact sequences (e.g. resolutions).
@@
    More generally, let $A$ be a regular local ring, then all finitely generated projective modules are free and hence they have a well-defined rank. Like in the case of fields, and in view of Proposition 2.9, we deduce that $K_{0}\left(A\right)\simeq\mathbb{Z}$ induced by the rank function.
Exercise 2.11.
@@
defines a commutative ring structure on $K_{0}\left(R\right)$. Make sure to spot where projectivity is used.
@@
makes $G_{0}\left(R\right)$ into a commutative ring. Moreover the inclusion $K_{0}\left(R\right)\to G_{0}\left(R\right)$ is a ring isomorphism (cfr. Proposition 2.9 and Exercise 2.11).
@@
In view of the explicit definition of the inverse map
$G_{0}\left(R\right)\to K_{0}\left(R\right),$
we see that requiring it to be a ring isomorphism amounts to the following: given two finitely generated $R$-modules $M,N$, choose two finite resolutions
$0\to P_{n}\to\dots\to P_{1}\to P_{0}\to M\to 0$
$0\to Q_{m}\to\dots\to Q_{1}\to Q_{0}\to N\to 0$
made up of finitely generated projective $R$-modules (they exist because the ring $R$ is supposed to be regular). Then we need to show that
$\sum_{i,j\geq 0}\left(-1\right)^{i+j}\left[P_{i}\otimes Q_{j}\right]=\sum_{k\geq 0}\left(-1\right)^{k}\left[\mathrm{Tor}_{k}^{R}\left(M,N\right)\right].$
Consider the bounded double complex $C_{\bullet.\bullet}$ with
$C_{i,j}=P_{i}\otimes Q_{j}.$
Then, in view of [Rot, Theorem 10.16, page 629], both spectral sequences associated to $C_{\bullet,\bullet}$ converge to the homology of $\mathrm{Tot}\left(C_{\bullet,\bullet}\right)_{\bullet}$ which is [^8] In fact, as th modules $P_{i}$’s and $Q_{j}$’s are all projective, all lines and columns are exact, therefore at the next page we find that the only non-zero row (or column) is given by this sequence of $\mathrm{Tor}$’s.
$H_{n}\left(\mathrm{Tot}\left(C_{\bullet,\bullet}\right)\right)=\mathrm{Tor}_{n}^{R}\left(M,N\right).$
In view of Lemma 2.8 we deduce our desidered equality. ∎
@@
I leave as an exercise to the willing reader to show directly that this definition really endows $G_{0}\left(R\right)$ with a ring structure. The only non trivial point to show is associativity, for which I know no proof (maybe using a spectral sequence for nested $\mathrm{Tor}$’s?).
@@
Let $\mathcal{F}:\mathcal{C}\to\mathcal{D}$ be a left exact additive functor between two abelian categories and suppose that its right derived functors $\left\{R^{i}\mathcal{F}\right\}_{i\geq 0}$ exist and that, for every object $X$, we have $R^{i}\mathcal{F}\left(X\right)=0$ for $i\gg 0$ (depending on $X$). Then it is well defined a group homomorphism
@@
Our main example will be the following: let $f:X\to Y$ be a morphism between projective varieties, then we define $f_{!}:K_{0}\left(X\right)\to K_{0}\left(Y\right)$ as $\left(f_{{}_{\ast}}\right)_{!}$. Note that $f$ has finite relative dimension, so that our assumptions are satisfied.
@@
Let $f:X\to Y$ be a closed embedding between two projective varieties with $Y$ smooth. For a vector bundle $\mathcal{F}$ on $X$, let
$0\to\mathcal{F}_{r}\to\mathcal{F}_{r-1}\to\dots\to\mathcal{F}_{0}\to f_{\ast}\mathcal{F}\to 0$
be a locally free coherent resolution of $f_{\ast}\mathcal{F}$. Then the assignment
$f_{!}:K_{0}\left(X\right)$ $\to K_{0}\left(Y\right)$
is given by
$f_{!}\left[\mathcal{F}\right]=\sum_{i=0}^{r}\left(-1\right)^{i}\left[\mathcal{F}_{i}\right].$
@@
First, let us note that $R^{i}f_{\ast}=0$ for $i>0$ and then, as our varieties are projective, in view of [Har, Theorem III.8.8, b), pag. 252] the functor $f_{\ast}\left(\bullet\right)$ preserves coherence and we have a well-defined map
$G_{0}\left(X\right)$ $\to G_{0}\left(Y\right)$
$\left[M\right]$ $\mapsto\left[f_{\ast}M\right].$
Then $f_{!}$ is just the composition
$K_{0}\left(X\right)\hookrightarrow G_{0}\left(X\right)\to G_{0}\left(Y\right)\simeq K_{0}\left(Y\right),$
where the last isomorphism is the one in Proposition 2.9. ∎
@@ 2.2 An important example: the Grothendieck group of projective spaces
$\phi\left(x^{t}\right)=\left[\mathcal{O}_{X}\left(t\right)\right].$
@@
$0\to\mathcal{O}_{X}\to\bigoplus_{i=1}^{n+1}\mathcal{O}_{X}\left(1\right)\to\bigoplus_{i=1}^{{n+1\choose 2}}\mathcal{O}_{X}\left(2\right)\to\dots\to\mathcal{O}_{X}\left(n+1\right)\to 0.$
In view of Lemma 2.7, denoting with $\xi=\left[\mathcal{O}_{X}\left(1\right)\right]$, we have
@@
which shows in turn that $\phi$ factors thought
$\tilde{\phi}:\frac{\mathbb{Z}\left[x\right]}{\left(1-x\right)^{n+1}}\to K_{0}\left(X\right)$
@@
$\left[\mathcal{O}_{X}\left(-1\right)\right]=\phi\left(\left(-1\right)^{n+1}\xi^{n}+\dots+{n+1\choose 2}\xi-\left(n+1\right)\right).$
In particular $\tilde{\phi}$ is surjective and $K_{0}\left(X\right)$ is generated by the classes
$\left[\mathcal{O}_{X}\right],\left[\mathcal{O}_{X}\left(1\right)\right],\dots,\left[\mathcal{O}_{X}\left(n\right)\right],$
that is, by $1,\xi,\dots,\xi^{n}$. Let
@@
$0=\alpha\cdot\xi^{-j}=a_{0}\left[\mathcal{O}_{X}\left(-j\right)\right]+\dots+a_{j}\left[\mathcal{O}_{X}\right],$
@@
$0=\chi\left(X,\alpha\cdot\xi^{-j}\right)=a_{j}\cdot\chi\left(X,\mathcal{O}_{X}\right)=a_{j}$
@@ 3.1 The Chow groups
$\left(f\right)=\sum_{V\subseteq W}\mathrm{ord}_{V}\left(r\right)\cdot\left[V\right]$
where the sum is take over codimension one subvarieties $V$ of $W$ and $\mathrm{ord}_{V}$ is the order function associated to the 1-dimensional local ring $\mathcal{O}_{W,V}$ [^10] You may want to glance at [Har, Exercise I.3.13, page 22] (see Lemma 3.1). We say that a $k$-cycle $\alpha$ is rationally equivalent to 0, written $\alpha\sim 0$, if
@@
Let $X$ be irreducible and let $f,g$ be finite morphisms, then $\deg\left(f\circ g\right)=\deg\left(f\right)\circ\deg\left(g\right)$.
@@
3.1.1 Intersection with divisors
In this section, in view of what follows, we have to deal with possibly singular varieties, referring to [Ful] for most of the proofs.
@@ 3.1.1 Intersection with divisors
Note that $D$ appears in the form of $\mathcal{O}_{X}\left(D\right)$, therefore we get a pairing
@@
One pleasant feature of the pairing just introduced comes from the following Proposition
@@ 3.2 Projective bundles
On $\mathbb{P}_{X}\left(\mathcal{E}\right)$ there is another “distinguished” sheaf, namely the pull-back $\pi^{\ast}\mathcal{E}$: it is locally free of rank $r+1$ (since $\mathcal{E}$ is) and a choice of a local basis amounts to the choice of local coordinates for $\mathbb{P}_{X}\left(\mathcal{E}\right)$, hence we have a natural map $\pi^{\ast}\mathcal{E}\to\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ which is surjective as $\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ is generated by the coordinates.
@@
$\alpha$ $\mapsto\pi_{\ast}\left(\left(r+i\right)D_{\mathcal{E}}\cdot\pi^{\ast}\alpha\right),$
@@
    $\alpha$ $\mapsto\left(-1\right)^{i}iD\cdot\alpha.$
@@
Theorem 3.16 ((Splitting construction)).
@@
    of locally free $\mathcal{O}_{Y}$-modules such that $\frac{\mathcal{E}_{i}}{\mathcal{E}_{i-1}}$ is locally free of rank 1 for every $i=1,\dots,r$.
@@
moreover, since $\mathrm{rk}\mathcal{O}_{Y^{\prime}}\left(1\right)=1$, we have
$\mathrm{rk}\tilde{\mathcal{E}}$ $=\mathrm{rk}\left(f^{\prime}\right)^{\ast}\mathcal{E}-1$
$=\mathrm{rk}\mathcal{E}-1$
$=r,$
we conclude by induction and Corollary 3.15. ∎

### v2

@@ 2 The $K$-theory
Summary
For the reader interested primarily in the proof of the Hirzebruch-Riemann-Roch Theorem, the technical construction of K-theory can be treated as a toolbox whose main properties we now outline.
To every smooth projective variety $X$, we associate an abelian group $K_{0}(X)$ generated by isomorphism classes of locally free sheaves (vector bundles). Since $X$ is smooth, any coherent sheaf admits a finite resolution by vector bundles; this fundamental result allows us to identify $K_{0}(X)$ with the group $G_{0}(X)$ generated by all coherent sheaves (Theorem 2.8). This group carries a commutative ring structure induced by the tensor product: $\mathcal{F}]\cdot[\mathcal{G}]=[\mathcal{F}\otimes\mathcal{G}]$ (Remark 2.10).
A crucial feature of the theory is its functoriality: any projective morphism $f:X\to Y$ induces a pushforward homomorphism $f_{!}:K_{0}(X)\to K_{0}(Y)$, defined via the alternating sum of higher direct images (Definition 2.14). This operation recovers classical invariants in specific cases: if $f$ is the map to a point, $f_{!}[\mathcal{F}]$ is the Euler characteristic $\chi(X,\mathcal{F})$ (Example 2.16); if $i:Z\hookrightarrow X$ is a closed embedding, $i_{!}$ simplifies to the classical direct image $i_{*}$ (Proposition 2.17).
Finally, explicitly computing these groups is possible: for the projective space $\mathbb{P}^{n}$, the ring $K_{0}(\mathbb{P}^{n})$ is
$K_{0}(\mathbb{P}^{n})\simeq\frac{\mathbb{Z}[x]}{(1-x)^{n+1}}$
where $x$ corresponds to $\mathcal{O}_{\mathbb{P}^{n}}(1)]$ (Theorem 2.19).
@@ 2.1 The Grothendieck group
This is a particular case of Example 2.13 since an exact functor has vanishing higher derived functors ($R^{i}\mathcal{F}=0$ for $i>0$).
2.2 Dévissage theorem
We see here how, in case the base variety or ring is regular, the $K$-groups we get considering coherent sheaves are the same that we get considering only locally free coherent sheaves.
@@ 2.2 Dévissage theorem
Let $\mathcal{P}$ be a category with exact sequences seen as a full subcategory of an abelian category $\mathcal{C}$ and suppose that, for every epimorphism $\pi$ in $\mathcal{P}$, $\ker\left(\pi\right)$ is still an object of $\mathcal{P}$. Let
@@
be an exact sequence in $\mathcal{P}$. Then the equality
@@
Let us proceed by induction on $n$: the cases when $n\leq 2$ are clear. Suppose the statement holds for n1n-1, then we can consider the two exact sequences
@@
Theorem 2.8.
[Dévissage] Let $A$ be a regular ring [^6] We say that a local ring $\left(A,\mathfrak{m},k\right)$ is a regular local ring if it is noetherian and $\dim A=\dim_{k}\frac{\mathfrak{m}}{\mathfrak{m}^{2}}$ We say that a (non necessarily local) ring $A$ is regular if it is noetherian and for every prime ideal $\mathfrak{p}\in\mathrm{Spec}\left(A\right)$, the local ring $A_{\mathfrak{p}}$ is regular. (resp. let $X$ be a smooth projective variety). Then the natural inclusions
$K_{0}\left(A\right)\to G_{0}\left(A\right)\quad\quad\mbox{and}\quad\quad K_{0}\left(X\right)\to G_{0}\left(X\right)$
are isomorphisms. [^7] Look at the Resolution Theorem [Ros, Theorem 3.1.13, pag. 121] (or the subsequent page for the $K_{1}$-analogue) for a generalization of this result. In general one has a pletora of dévissage theorems for reducing the study of a category to that of a well-behaved subcategory by means of suitable exact sequences (e.g. resolutions).
@@
    More generally, let $A$ be a regular local ring, then all finitely generated projective modules are free and hence they have a well-defined rank. Like in the case of fields, and in view of Theorem 2.8, we deduce that $K_{0}\left(A\right)\simeq\mathbb{Z}$ induced by the rank function.
2.3 Multiplicative structure
Remark 2.10.
@@ 2.3 Multiplicative structure
defines a commutative ring structure on $K_{0}\left(R\right)$. More generally if $X$ is a projective variety then $K_{0}(X)$ becomes a commutative ring via
$\mathcal{F}]\cdot[\mathcal{G}]=[\mathcal{F}\otimes\mathcal{G}].$
If fact the definitions are well posed since locally free modules are flat, moreover associativity boils down to usual associativity for the tensor product.
Diving deeper into homological algebra one could prove
@@
makes $G_{0}\left(R\right)$ into a commutative ring such that the inclusion $K_{0}\left(R\right)\to G_{0}\left(R\right)$ is a ring isomorphism (cfr. Theorem 2.8 and Remark 2.10).
@@
This is the formula obtained transporting the ring structure of $K_{0}(R)$ via the Dévissage isomorphism of Theorem 2.8. See [Ser, Chapter V] for a discussion of intersection multiplicities via Tor. ∎
@@
I leave as an exercise to the willing reader to show directly that this definition really endows $G_{0}\left(R\right)$ with a ring structure. The only non trivial point to show is associativity.
2.4 Functoriality
@@
Let $\mathcal{F}:\mathcal{C}\to\mathcal{D}$ be a left exact additive functor between two abelian categories and suppose that its right derived functors $\left\{R^{i}\mathcal{F}\right\}_{i\geq 0}$ exist and that, for every object $X$, we have $R^{i}\mathcal{F}\left(X\right)=0$ for $i\gg 0$ (depending on $X$). Then the assignment
@@
is a well defined group homomorphism: if $0\to X\to Y\to Z\to 0$ is an exact sequence in $\mathcal{C}$ then we have an exact sequence
$0\to\mathcal{F}(X)\to\mathcal{F}(Y)\to\mathcal{F}(Z)\to R^{1}\mathcal{F}(X)\to R^{1}\mathcal{F}(Y)\to R^{1}\mathcal{F}(Z)\to\dots$
that is finite by assumption. Since $\mathcal{D}$ is abelian it has kernels hence Lemma 2.6 tells that
$\mathcal{F}_{!}\left[Y\right]=\mathcal{F}_{!}\left[X\right]+\mathcal{F}_{!}\left[Z\right]$
Our main example will be the following: let $f:X\to Y$ be a morphism between projective varieties, in particular it has finite relative dimension and $R^{i}f_{\ast}$ preserves coherence so the discussion above applies to $f_{\ast}$ hence $(f_{\ast})_{!}:G_{0}\left(X\right)\to G_{0}\left(Y\right)$ is well defined and, when $Y$ is smooth, it restricts to $K_{0}(X)\to K_{0}(Y)$ in view of Theorem 2.8.
Definition 2.14.
Let $f:X\to Y$ be a morphism between projective varieties, we denote
$f_{!}=(f_{\ast})_{!}:G_{0}\left(X\right)\to G_{0}\left(Y\right)$
the map introduced in Example 2.13. When $Y$ is smooth with an abuse of notation we also denote
$f_{!}:K_{0}\left(X\right)\to K_{0}\left(Y\right)$
the induced map.
@@
We give now some examples
Example 2.16.
Let $X$ be a projective variety and consider the map $f:X\to\{\ast\}$ to a point. Clearly the category of coherent sheaves on a point is just the category of finite dimensional vector spaces and they are classified by their dimension, hence
$K_{0}(\{\ast\})=G_{0}(\{\ast\})=\mathbb{Z}.$
Since $R^{i}f_{\ast}\mathcal{F}=H^{i}\left(X,\mathcal{F}\right)$ we have $f_{!}\mathcal{F}=\sum_{i\geq 0}(-1)^{i}\dim_{k}H^{i}\left(X,\mathcal{F}\right)$ that is
$f_{!}\mathcal{F}=\chi\left(X,\mathcal{F}\right)$
is the Euler characteristic of $\mathcal{F}$.
@@
Let $f:X\to Y$ be a closed embedding between two projective varieties with $Y$ smooth. If $\mathcal{F}$ is a coherent sheaf on $X$ then
  1. 1.
    $f_{!}[\mathcal{F}]=[f_{\ast}\mathcal{F}]$
  2. 2.
    $\chi\left(Y,j_{!}[\mathcal{F}]\right)=\chi\left(X,[\mathcal{F}]\right)$
@@
If $\mathcal{F}$ is a coherent sheaf on $X$ then, since $R^{i}f_{\ast}\mathcal{F}$ is the sheaf associated with the presheaf
$U\mapsto H^{i}\left(f^{-1}\left(U\right),\mathcal{F}_{|f^{-1}\left(U\right)}\right),$
we see that $R^{i}f_{\ast}\mathcal{F}=0$ for $i\geq 1$ and $f_{\ast}\mathcal{F}$ is just extension by 0 outside the image of $f$ (see [Har, Proposition III.8.1, pag. 250]). Then
$f_{!}[\mathcal{F}]=[f_{\ast}\mathcal{F}]$
for every coherent sheaf $\mathcal{F}$ on $X$. In particular we compute:
$\chi\left(Y,j_{!}[\mathcal{F}]\right)=\sum_{i\geq 0}\left(-1\right)^{i}\chi\left(Y,\left[R^{i}j_{\ast}\mathcal{F}\right]\right)=\chi\left(Y,\left[j_{\ast}\mathcal{F}\right]\right)=\chi\left(X,\left[\mathcal{F}\right]\right).$
@@ 2.5 An important example: the Grothendieck group of projective spaces
[…]
@@
which shows in turn that $\phi$ factors through
[…]
@@
[…]
that is, by $1,\xi,\dots,\xi^{n}$. To conclude, we show that $1,\xi,\dots,\xi^{n}$ are linearly independent over $\mathbb{Z}$. Let
@@
[…]
@@
[…]
@@ 3.1 The Chow groups
$\left(f\right)=\sum_{V\subseteq W}\mathrm{ord}_{V}\left(f\right)\cdot\left[V\right]$
where the sum is taken over codimension one subvarieties $V$ of $W$ and $\mathrm{ord}_{V}$ is the order function associated to the 1-dimensional local ring $\mathcal{O}_{W,V}$ [^9] You may want to glance at [Har, Exercise I.3.13, page 22] (see Lemma 3.1). We say that a $k$-cycle $\alpha$ is rationally equivalent to 0, written $\alpha\sim 0$, if
@@
Let $X$ be irreducible and let $f,g$ be finite morphisms, then
$\deg\left(f\circ g\right)=\deg\left(f\right)\circ\deg\left(g\right)$
@@
3.2 Intersection with divisors
@@
In view of the smoothness assumption the pairing in Definition 3.7 induces
@@ 3.3 Projective bundles
On $\mathbb{P}_{X}\left(\mathcal{E}\right)$ there is another “distinguished” sheaf, namely the pull-back $\pi^{\ast}\mathcal{E}$: it is locally free of rank $r+1$ (since $\mathcal{E}$ is) and a choice of a local basis amounts to the choice of local coordinates for $\mathbb{P}_{X}\left(\mathcal{E}\right)$, hence we have a natural map $\pi^{\ast}\mathcal{E}\to\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ which is surjective as $\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ is generated by its sections.
@@
$\alpha$ $\mapsto\pi_{\ast}\left(D^{r+i}_{\mathcal{E}}\cdot\pi^{\ast}\alpha\right),$
@@
    $\alpha$ $\mapsto\left(-1\right)^{i}D^{i}\cdot\alpha.$
@@
Theorem 3.16 (Splitting construction).
@@
    of locally free $\mathcal{O}_{Y}$-modules such that $\dfrac{\mathcal{E}_{i}}{\mathcal{E}_{i-1}}$ is locally free of rank 1 for every $i=1,\dots,r$.
@@
moreover, since $\mathrm{rk}\left(\mathcal{O}_{Y^{\prime}}\left(1\right)\right)=1$, we have $\mathrm{rk}\left(\tilde{\mathcal{E}}\right)=\mathrm{rk}\left(\left(f^{\prime}\right)^{\ast}\mathcal{E}\right)-1=\mathrm{rk}\left(\mathcal{E}\right)-1=r$ we conclude by induction and Corollary 3.15. ∎